Circular No. 5917

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRO J0422+32 E. L. Martin and M. Serra-Ricart, Instituto de Astrofisica de Canarias,

report: “The x-ray transient GRO J0422+32, a possible black-hole binary, is currently returning to its quiescent state. CCD observations at the Teide Observatory 0.80-m IAC telescope give the following magnitudes in the R band: 1993 Nov. 27.25 UT, 18.5; Dec. 14.02, 14.9; 15.97, 15.0; 17.03, 15.0; 18.12, 15.2; 1994 Jan. 4.11, 18.0; 4.98, 18.9. The last two nights of data show that the object has rapidly faded by more than 3 mag and indicate that this outburst (IAUC 5901) has been short-lived (duration < 40 days, as with that of August 1993; IAUC 5842).”

SUPERNOVA 1993aj IN ANONYMOUS GALAXY A. V. Filippenko and T. Matheson, University of California at Berke-

ley, report that an uncalibrated CCD spectrogram (range 310-1000 nm) obtained on Jan. 7 UT with the 3-m Shane reflector at Lick Observatory suggests that this object (IAUC 5915) is a type-II supernova, within a few weeks of maximum brightness. Broad hydrogen Balmer lines with P-Cyg profiles are visible.

PERIODIC COMET MUELLER 5 (1993s)

Improved orbital elements from MPEC 1994-A04:

T = 1994 Sept.19.763 TT  ω = 31.^{o}023
e = 0.26243  Ω = 100.658 }2000.0
q = 4.24259 AU  i = 16.491

a = 5.75214 AU n^{o} = 0.071443 P = 13.80 years

1994 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Jan. 8 6^{h}34.^{m}53 +22^{o}59.^{′}2 3.405 4.378 170.^{o}4 2.^{o}2 17.6
18 6 28.67 +23 29.6 3.436 4.368 158.8  4.7 17.6
28 6 23.69 +23 58.0 3.495 4.358 147.5  7.0 17.6
Feb. 7 6 20.02 +24 23.8 3.580 4.349 136.5  9.0 17.7
17 6 17.92 +24 47.0 3.686 4.340 125.9 10.6 17.7
27 6 17.55 +25 07.7 3.808 4.331 115.7 11.9 17.8
Mar. 9 6 18.91 +25 26.0 3.942 4.323 106.0 12.8 17.8
19 6 21.96 +25 41.9 4.084 4.315  96.7 13.2 17.9
29 6 26.58 +25 55.3 4.228 4.307  87.8 13.4 18.0

1994 January 7 Daniel W. E. Green